Sixteen ceramic tiles are needed to tile a 9-square-foot area. At this rate, how many square feet can be tiled 256 ceramic tiles?
Answer by stanbon(75887)

Sixteen ceramic tiles are needed to tile a 9-square-foot area. At this rate, how many square feet can be tiled 256 ceramic tiles?
---
x/256 = 9/16
----
x = (256/16)*9 =
---
x = 144 sq ft
===============
Cheers,
Stan H.
===============